# SPAGS: Sparse-View Articulated Object Reconstruction from Single State via Planar Gaussian Splatting

### v2

@@ Abstract
Articulated objects are ubiquitous in daily environments, and their 3D reconstruction holds great significance across various fields. However, existing articulated object reconstruction methods typically require costly inputs such as multi-stage and multi-view observations. To address the limitations, we propose a category-agnostic articulated object reconstruction framework via planar Gaussian Splatting, which only uses sparse-view RGB images from a single state. Specifically, we first introduce a Gaussian information field to perceive the optimal sparse viewpoints from candidate camera poses. Then we compress 3D Gaussians into planar Gaussians to facilitate accurate estimation of normal and depth. The planar Gaussians are optimized in a coarse-to-fine manner through depth smooth regularization and few-shot diffusion. Moreover, we introduce a part segmentation probability for each Gaussian primitive and update them by back-projecting part segmentation masks of renderings. Extensive experimental results demonstrate that our method achieves higher-fidelity part-level surface reconstruction on both synthetic and real-world data than existing methods. Codes will be made publicly available.
1 Introduction
@@
Under this circumstance, some methods [mu2021learning, jiang2022ditto] learn category-specific prior knowledge for articulated object reconstruction. Nevertheless, these models struggle with unseen object types. Recently, PARIS [jiayi2023paris], REArtGS [wu2025reartgs], and ArtGS [liu2025building] achieve category-agnostic part-level reconstruction using multi-view RGB or RGBD images from two states of articulated objects. However, the multi-view observations from two states are costly in certain cases, especially for robot manipulation, which requires robots to perform time-consuming motion and complex pose estimation for image acquisition.
@@
In recent years, some approaches [zhu2023FSGS, paliwal2024coherentgs] are dedicated to reducing reliance on multi-view reconstruction. Unfortunately, when the input views become extremely sparse (e.g., 4 views), their reconstruction quality suffers a sharp decline. Most recently, GaussianObject [yang2024gaussianobject] incorporates diffusion model [song2020denoising] with 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians], achieving realistic object reconstruction only with 4-view RGB images. Nevertheless, we make observation that GaussianObject still exhibits followings limitations for articulated object reconstruction: (1) The disorderly and irregular nature of 3D Gaussian primitives makes depth and normal estimation difficult, leading to inaccurate surface reconstruction. (2) It requires a predefined view setting, and there is a deviation between the manually selected viewpoints and the optimal observation, affecting its application in robot autonomous perception. (3) It fails to extract part-level meshes, limiting downstream manipulation of articulated objects.
@@
To tackle the above-mentioned challenges, we propose SPAGS (shown in Fig. 1), the first category-agnostic framework of single-state SParse-view Articulated object reconstruction via Gaussian Splatting, to our best knowledge. Specifically, we first propose an optimal view perception method by establishing a Gaussian information field, which continuously estimates optimal viewpoints with the maximum information potential. We later acquire a structured initialization for Gaussian primitives via a 3D generative model [huang2025spar3d] and register the initialization using a pyramid network.
@@
Afterwards, we compress 3D Gaussians into planar Gaussians to facilitate depth and normal estimation. We adopt a coarse-to-fine strategy to optimize the planar Gaussians. In coarse training, we leverage the pseudo depth labels from a depth estimation model and depth smoothness regularization to enhance geometric learning. During refinement, we fine-tune a pretrained diffusion model using few-shot data by constructing image pairs within reliable regions, and then use the fine-tuned model to refine noisy regions. Subsequently, we employ a Visual Language Model (VLM) to generate an articulation tree, and assign part-aware probability to each Gaussian primitive, which is updated via back-projecting 2D part masks. Extensive experiments demonstrate our method significantly outperforms the existing state-of-the-art (SOTA) methods on both synthetic and real-world data.
In summary, our main contributions can be summarized as follows: (1) We propose the first category-agnostic framework of high-fidelity articulated object reconstruction, using only sparse-view RGB images from a single state. (2) We propose the Gaussian information field to estimate the information potential of a sampled camera pose, achieving optimal view perception. (3) We propose a coarse-to-fine optimization strategy for planar Gaussians, improving geometry learning through depth smooth regularization and few-shot diffusion.
@@
Figure 2: The Framework of SPAGS. We use the snowflake symbol to denote frozen network weights and the flame symbol to indicate trainable weights. “Reg.” and “Regist.” denote regularization and registration respectively. We highlight our main contributions in green. Our method SPAGS can autonomously perceive the optimal sparse viewpoints and achieve high-fidelity reconstruction results for arbitrary articulated objects.
2 Related Works
2.1 Sparse-View Image Reconstruction
@@
While NeRF [mildenhall2020nerf] and 3DGS [kerbl3Dgaussians] emerge as novel fashions for high-fidelity scene reconstruction, these methods typically struggle with sparse-view input setting. Therefore, some works attempt to reduce the number of training views. Specifically, RegNeRF [Niemeyer2021Regnerf] proposes additional geometric regularization from unobserved viewpoints. SparseNeRF [wang2022sparsenerf] distills local depth ranking priors from a monocular depth estimation model. Although achieving improved novel view synthesis results, they yield entangled representation from the implicit radiation fields, making it challenging to extract an acceptable surface mesh. Recently, FSGS [zhu2023FSGS] introduces a proximity-guided Gaussian unpooling to increase the density of 3D Gaussians with sparse-view training images. CoherentGS [paliwal2024coherentgs] proposes a coherent regularization to optimize structured 3D Gaussians. Sparse2DGS [Wu_2025_CVPR] combines MVS with 3DGS for improved sparse-view reconstruction. SparseGS [xiong2023sparsegs] and GaussianObject [yang2024gaussianobject] integrate the diffusion model with 3DGS for sparse 360-degree reconstruction. However, these methods still rely on pre-defined sparse-view selection, and the irregular nature of 3D Gaussian primitives hinders accurate normal and depth estimation.
2.2 Articulated Object Surface Reconstruction
@@
Articulated object surface reconstruction has gradually attracted more attention in recent years. ASDF [mu2021learning] and Ditto [jiang2022ditto] learn pre-trained models for shape generation of articulated objects, but they struggle with unseen object types. SINGAPO [liu2024singapo] and DreamArt [lu2025dreamart] propose single-view surface generation methods for articulated objects. However, these methods usually yield unfaithful results in unseen regions. Recently, PARIS [jiayi2023paris], ArticulatedGS [guo2025articulatedgs], REArtGS [wu2025reartgs] and ArtGS [liu2025building] achieves category-agnostic high-fidelity reconstruction of articulated objects with two-stage multi-view RGB images. However, these methods still require costly inputs, limiting their practicality in downstream tasks.
3 Methodology
@@
3.1 Optimal Viewpoint Perception and Representation Initialization
@@
where $\alpha$ is the $\alpha$-blending weight and $P(1:k)\in\left[0,1\right]$ is current reliability, quantifying rendering quality via LPIPS [zhang2018perceptual] and NIQE [NIQE], without ground truth image. Here, we extend LPIPS result to a pixel-wise heatmap obtained by upsampling the spatial error map. We define the information potential $E$ as: $E=-\log P(1:k)$. Intuitively, $\Psi$ encodes Gaussian reliable probability, forming a ray-modulated energy field.
@@
where $n$ means all measurement beams, and $N_{\mathcal{G}}$ denotes the number of the ordered Gaussians along the ray. We incorporate the $k+1$ optimal viewpoint $\sigma^{*}$ into existing viewpoint set $\pi$, and utilize the captured images from updated $\pi$ to optimize the Gaussian primitives by vanilla 3DGS pipeline [kerbl3Dgaussians]. Besides, we update $P$ through the accumulated information potential $E$ to ensure $P$ within the defined bounds $[0,1]$. The accumulated information potential $E(1:k+1)$ is formulated as:
@@
Note that GauSS-MI [xie2025gaussmi] proposes a method similar to ours, but GauSS-MI requires the observations or expected RGB results corresponding to the candidate viewpoints. In contrast, our method conducts optimal viewpoint perception only with candidate camera poses from random sampling, which is more convenient for downstream applications.
@@
Once we finish sparse view perception, we attempt to obtain an initial initialization $\mathbf{p}\in\mathbb{R}^{3}$ using a monocular 3D generation model SPAR3D [huang2025spar3d]. However, this is not a ready-to-use pipeline since the generated pose and scale are inaccurate, which seriously affects subsequent reconstruction. We observe that although the points $\mathbf{\bar{p}}\in\mathbb{R}^{3}$ of Gaussian primitives from view perception are noisy, the global pose and scale of $\mathbf{\bar{p}}$ are relatively accurate. Therefore, we seek to distill a structured representation with reasonable pose and scale through registering $\mathbf{p}$ and $\mathbf{\bar{p}}$.
Due to the complex disparities in pose and scale between $\mathbf{p}$ and $\mathbf{\bar{p}}$, we assume that each individual point $\mathbf{p}_{i}\in\mathbf{p}$ undergoes a similarity transformation to $\mathbf{\hat{p}}_{i}$. We employ an exponential mapping $\mathbf{T}=\left(\mathbf{\omega},\mathbf{t},s\right)$ to parameterize the similarity transformation, where $s$ is the scaling factor and $\mathbf{t}$ is the translation. $\mathbf{\omega}\in\mathbb{R}^{3}$ is an axis-angle vector. Inspired by [li2022DeformationPyramid], we adopt a pyramid structure network to progressively learn the deformation parameters $\mathbf{T}$ which is decomposed into a sequence $\left\{\mathbf{T}_{1},\mathbf{T}_{2},...,\mathbf{T}_{l}\right\}$, with $l=9$. Please refer to the supplementary material for detailed elaboration.
3.2 Coarse Training for Planar 3D Gaussians
@@
where $S_{1},S_{2},S_{3}$ denote the scale of Gaussian primitive $\mathcal{G}_{i}$ along each direction. Based on the planar Gaussian primitives, we can easily acquire the normal via the shortest axis and viewing direction. Denoting the normal rendering from $\alpha$-blending as NN, we can derive the unbiased depth $\mathbf{D}$ following PGSR [chen2024pgsr], which are illustrated in the supplementary materials.
Table 1: Quantitative results for the surface reconstruction quality on PartNet-Mobility dataset. denotes methods designed for multi-view reconstruction. We bold the best results and underline the second best results.
CD $\downarrow$ PGSR [chen2024pgsr] 13.19 26.94 6.01 672.76 32.20 32.40 2.49 24.12 48.68 81.90 94.07
CoherentGS [paliwal2024coherentgs] 324.65 33.76 313.04 40.87 80.61 64.31 9.34 164.08 41.42 303.67 137.58
Sparse2DGS [Wu_2025_CVPR] 49.25 35.57 28.96 38.87 12.64 425.27 4.13 28.47 73.74 60.88 75.78
GaussianObject [yang2024gaussianobject] 7.39 6.92 0.82 36.05 0.58 68.87 1.58 166.03 12.30 126.93 42.75
SPAGS (Ours) 5.65 6.56 0.89 9.04 0.66 27.54 1.05 5.21 17.56 21.80 9.60
F1 $\uparrow$ PGSR [chen2024pgsr] 0.08 0.13 0.33 0.02 0.15 0.05 0.25 0.19 0.03 0.01 0.12
CoherentGS [paliwal2024coherentgs] 0.01 0.06 0.08 0.06 0.15 0.03 0.32 0.17 0.02 0.00 0.09
Sparse2DGS [Wu_2025_CVPR] 0.11 0.07 0.24 0.05 0.18 0.03 0.16 0.13 0.03 0.04 0.10
GaussianObject [yang2024gaussianobject] 0.10 0.15 0.38 0.05 0.35 0.05 0.27 0.10 0.03 0.01 0.15
SPAGS (Ours) 0.14 0.19 0.40 0.06 0.38 0.05 0.32 0.13 0.03 0.02 0.17
EMD $\downarrow$ PGSR [chen2024pgsr] 0.14 0.19 0.07 0.84 0.16 0.17 0.06 0.13 0.22 0.23 0.22
CoherentGS [paliwal2024coherentgs] 0.53 0.26 0.44 0.20 0.32 0.23 0.06 0.16 0.25 0.60 0.31
Sparse2DGS [Wu_2025_CVPR] 0.50 0.59 0.64 0.62 0.30 0.72 0.50 0.51 0.28 0.31 0.50
GaussianObject [yang2024gaussianobject] 0.17 0.12 0.09 0.12 0.09 0.20 0.05 0.31 0.17 0.38 0.17
SPAGS (Ours) 0.12 0.11 0.09 0.10 0.08 0.17 0.04 0.10 0.16 0.20 0.12
To enhance geometric constraints for the planar Gaussians, we employ a monocular depth estimation model [depthanything] to introduce additional depth cues $\mathcal{D}$ for sparse-view images. Considering that the depth estimation result $\mathcal{D}$ is typically noisy, we employ scale parameter $\boldsymbol{\varphi}=\left\{\varphi_{1},\varphi_{2},...,\varphi_{K}\right\}$ and offset parameter $\boldsymbol{\eta}=\left\{\eta_{1},\eta_{2},...,\eta_{K}\right\}$ to adaptively refine the estimated depth $\mathcal{D}_{i}$ for each view. Specifically, we set $\boldsymbol{\varphi}$ as a learnable parameter to align depth scale. We optimize the parameters of an image convolutional decoder to update the offset $\boldsymbol{\eta}$ of whole image instead of estimating pixel-wise offset, ensuring a coherent variation. Given the depth rendering $\mathbf{D}$ of planar Gaussians, the depth regularization $\mathcal{L}_{\text{depth}}$ can be formulated as:
@@
Moreover, we encourage the depth variation of Gaussian primitives to be smooth. We start with adopting the $c$-channel segmentation module from [paliwal2024coherentgs] to segment each estimated depth $\mathcal{D}_{i}$ into $c$ similar regions, as shown in part (3) of Fig. 2. For each region, we introduce a local depth constraint $\mathcal{L}_{\text{smooth}}$ to ensure the depth variations are smooth, which are expressed as:
@@
The coarse training objective can be formulated as: $\mathcal{L}_{\text{coarse}}=\mathcal{L}_{\text{color}}+\mathcal{L}_{\text{planar}}$ , where $\mathcal{L}_{\text{color}}$ is the color loss of pixel space using in [kerbl3Dgaussians]. $\mathcal{L}_{\text{planar}}$ is expressed as:
@@
3.3 Planar Gaussian Refinement via Few-shot Diffusion
To optimize the representation of regions lacking supervision, we introduce the diffusion model [rombach2022high] to futher refine Gaussian primitives. Note that our approach only utilize few-shot data to fine-tune the diffusion model, i.e., the sparse-view inputs.

### v4

@@ Abstract
Articulated objects are ubiquitous in daily environments, and their 3D reconstruction holds great significance across various fields. However, existing articulated object reconstruction methods typically require costly inputs such as multi-stage and multi-view observations. To address the limitations, we propose a category-agnostic articulated object reconstruction framework via planar Gaussian Splatting, which only uses sparse-view RGB images from a single state. Specifically, we first introduce a Gaussian information field to perceive the optimal sparse viewpoints from candidate camera poses. To ensure precise geometric fidelity, we constrain traditional 3D Gaussians into planar primitives, facilitating accurate normal and depth estimation. The planar Gaussians are then optimized in a coarse-to-fine manner, regularized by depth smoothness and few-shot diffusion priors. Furthermore, we leverage a Vision-Language Model (VLM) via visual prompting to achieve open-vocabulary part segmentation and joint parameter estimation. Extensive experiments on both synthetic and real-world datasets demonstrate that our approach significantly outperforms existing baselines, achieving superior part-level surface reconstruction fidelity. Code and data are provided in the supplementary material.
I Introduction
@@
Under this circumstance, some methods [23, 10] learn category-specific prior knowledge for articulated object reconstruction. Nevertheless, these models struggle with unseen object types. Recently, PARIS [16], REArtGS [33], and ArtGS [17] achieve category-agnostic part-level reconstruction using multi-view RGB or RGBD images from two states of articulated objects. However, obtaining multi-view observations from two states is costly in practical applications, especially in the task of active 3D reconstruction for robots, which relies on time-consuming robot movements and complex inverse kinematics computation to complete image acquisition, which is illustrated in Fig. 3.
@@
In recent years, some approaches [45, 26] are dedicated to reducing reliance on multi-view reconstruction. Unfortunately, when the input views become extremely sparse (e.g., 4 views), their reconstruction quality suffers a sharp decline. Most recently, GaussianObject [40] incorporates diffusion model [30] with 3D Gaussian Splatting (3DGS) [11], achieving realistic object reconstruction only with 4-view RGB images. Nevertheless, we make observation that GaussianObject still exhibits followings limitations for articulated object reconstruction: (1) The disorderly and irregular nature of 3D Gaussian primitives makes depth and normal estimation difficult, leading to inaccurate surface reconstruction. (2) It requires a predefined view setting, and there is a deviation between the manually selected viewpoints and the optimal observation, affecting its application in robot autonomous perception. (3) It fails to extract part-level meshes, limiting downstream manipulation of articulated objects.
@@
To tackle the above-mentioned challenges, we propose SPAGS (shown in Fig. 1), the first category-agnostic framework of single-state SParse-view Articulated object reconstruction via Gaussian Splatting, to our best knowledge. Specifically, we first propose an optimal view perception method by establishing a Gaussian information field, which continuously estimates optimal viewpoints with the maximum information potential. We later acquire a structured initialization for Gaussian primitives via a 3D generative model [8] and register the initialization using a pyramid network.
@@
Afterwards, we compress 3D Gaussians into planar Gaussians to facilitate depth and normal estimation. We adopt a coarse-to-fine strategy to optimize the planar Gaussians. In coarse training, we leverage the pseudo depth labels from a depth estimation model and depth smoothness regularization to enhance geometric learning. During refinement, we fine-tune a pretrained diffusion model using few-shot data by constructing image pairs within reliable regions, and then use the fine-tuned model to refine noisy regions. Subsequently, we employ a Visual Language Model (VLM) to generate an articulation tree, and assign part-aware probability to each Gaussian primitive, which is updated via back-projecting 2D part masks. Then we perform joint parameter estimation via visual question answering. Extensive experiments demonstrate our method significantly outperforms the existing state-of-the-art (SOTA) methods on both synthetic and real-world data.
In summary, our main contributions can be summarized as follows:
  • We propose the first category-agnostic framework of high-fidelity articulated object reconstruction, using only sparse-view RGB images from a single state.
  • We propose the Gaussian information field to estimate the information potential of a sampled camera pose, achieving optimal view perception.
  • We propose a coarse-to-fine optimization strategy for planar Gaussians, improving geometry learning through depth smooth regularization and few-shot diffusion.
@@
Figure 2: The Framework of SPAGS. We use the snowflake symbol to denote frozen network weights and the flame symbol to indicate trainable weights. “Reg.” and “Regist.” denote regularization and registration respectively. We highlight our main contributions in green. Note that the optimal view perception is an optional component.
II Related Works
II-A Sparse-View Image Reconstruction
@@
While 3DGS [11] and its subsequent works [36, 9, 7, 3, 14, 19, 29] emerge as novel fashions for high-fidelity scene reconstruction, these methods typically struggle with sparse-view input setting. Therefore, some works attempt to reduce the number of training views. Specifically, RegNeRF [24] proposes additional geometric regularization from unobserved viewpoints. SparseNeRF [31] distills local depth ranking priors from a monocular depth estimation model. Although achieving improved novel view synthesis results, they yield entangled representation from the implicit radiation fields, making it challenging to extract an acceptable surface mesh. Recently, FSGS [45] introduces a proximity-guided Gaussian unpooling to increase the density of 3D Gaussians with sparse-view training images. CoherentGS [26] proposes a coherent regularization to optimize structured 3D Gaussians. Sparse2DGS [34] combines MVS with 3DGS for improved sparse-view reconstruction. SparseGS [39] and GaussianObject [40] integrate the diffusion model with 3DGS for sparse 360-degree reconstruction. However, these methods still rely on pre-defined sparse-view selection, and the irregular nature of 3D Gaussian primitives hinders accurate normal and depth estimation. Recently, GauSS-MI [38] leverages RGB observations for iterative next-best-view estimation. However, this strategy requires image capture at every candidate pose, leading to a bottleneck in scenarios where image acquisition is infeasible. Therefore, achieving optimal viewpoint perception based solely on candidate camera poses offers significantly higher practical utility for real-world applications.
II-B Articulated Object Surface Reconstruction
@@
Articulated object surface reconstruction has gradually attracted more attention in recent years. ASDF [23] and Ditto [10] learn pre-trained models for shape generation of articulated objects, but they struggle with unseen object types. SINGAPO [15] and DreamArt [18] propose single-view surface generation methods for articulated objects. However, these methods usually yield unfaithful results in unseen regions. Recently, PARIS [16], ArticulatedGS [4], REArtGS [33], REArtGS++ [32] and ArtGS [17] achieves category-agnostic high-fidelity reconstruction of articulated objects with two-stage multi-view RGB images. However, these methods still require costly inputs, limiting their practicality in downstream tasks.
III Methodology
@@
III-A Optimal Viewpoint Perception and Representation Initialization
@@
where $\alpha$ is the $\alpha$-blending weight and $P(1:k)\in\left[0,1\right]$ is current reliability, quantifying rendering quality via NIQE [22] without ground truth image. We define the information potential $E$ as: $E=-\log P(1:k)$. Intuitively, $\Psi$ encodes Gaussian reliable probability, forming a ray-modulated energy field.
@@
where $n$ means all measurement beams, and $N_{\mathcal{G}}$ denotes the number of the ordered Gaussians along the ray. We incorporate the $k+1$ optimal viewpoint $\sigma^{*}$ into existing viewpoint set $\pi$, and utilize the captured images from updated $\pi$ to optimize the Gaussian primitives by vanilla 3DGS pipeline [11]. Besides, we update $P$ through the accumulated information potential $E$ to ensure $P$ within the defined bounds $[0,1]$. The accumulated information potential $E(1:k+1)$ is formulated as:
@@
Notably, the optimal viewpoint perception module is optional. It eliminates manual selection bias when viewpoint choice is feasible, yet our approach sustains high-quality 3D reconstruction results under the constraints of fixed and sparse viewpoints, as demonstrated in Table. VI.
@@
Figure 3: Comparison of the optimal sparse view acquisition and multi-view acquisition in robotic active reconstruction. Our method significantly reduces the complexity of the robot’s movement and the inverse kinematics computation.
Once we finish sparse view perception, we attempt to obtain an initial initialization $\mathbf{p}\in\mathbb{R}^{3}$ using a monocular 3D generation model SPAR3D [8]. However, this is not a ready-to-use pipeline since the generated pose and scale are inaccurate, which seriously affects subsequent reconstruction. We observe that although the points $\mathbf{\bar{p}}\in\mathbb{R}^{3}$ of Gaussian primitives from view perception are noisy, the global pose and scale of $\mathbf{\bar{p}}$ are relatively accurate. Therefore, we seek to distill a structured representation with reasonable pose and scale through registering $\mathbf{p}$ and $\mathbf{\bar{p}}$.
Due to the complex disparities in pose and scale between $\mathbf{p}$ and $\mathbf{\bar{p}}$, we assume that each individual point $\mathbf{p}_{i}\in\mathbf{p}$ undergoes a similarity transformation to $\mathbf{\hat{p}}_{i}$. We employ an exponential mapping $\mathbf{T}=\left(\mathbf{\omega},\mathbf{t},s\right)$ to parameterize the similarity transformation, where $s$ is the scaling factor and $\mathbf{t}$ is the translation. $\mathbf{\omega}\in\mathbb{R}^{3}$ is an axis-angle vector. Inspired by [13], we adopt a pyramid structure network to progressively learn the deformation parameters $\mathbf{T}$ which is decomposed into a sequence $\left\{\mathbf{T}_{1},\mathbf{T}_{2},...,\mathbf{T}_{l}\right\}$, with $l=9$. Please refer to the supplementary material for detailed elaboration.
III-B Coarse Training for Planar 3D Gaussians
@@
where $S_{1},S_{2},S_{3}$ denote the scale of Gaussian primitive $\mathcal{G}_{i}$ along each direction. Based on the planar Gaussian primitives, we can easily acquire the normal via the shortest axis and viewing direction. Denoting the normal rendering from $\alpha$-blending as $N$, we can derive the unbiased depth $\mathbf{D}$ following PGSR [1], which are illustrated in the supplementary materials.
To enhance geometric constraints for the planar Gaussians, we employ a monocular depth estimation model [41] to introduce additional depth cues $\mathcal{D}$ for sparse-view images. Considering that the depth estimation result $\mathcal{D}$ is typically noisy, we employ scale parameter $\boldsymbol{\varphi}=\left\{\varphi_{1},\varphi_{2},...,\varphi_{K}\right\}$ and offset parameter $\boldsymbol{\eta}=\left\{\eta_{1},\eta_{2},...,\eta_{K}\right\}$ to adaptively refine the estimated depth $\mathcal{D}_{i}$ for each view. Specifically, we set $\boldsymbol{\varphi}$ as a learnable parameter to align depth scale. We optimize the parameters of an image convolutional decoder to update the offset $\boldsymbol{\eta}$ of whole image instead of estimating pixel-wise offset, ensuring a coherent variation. Given the depth rendering $\mathbf{D}$ of planar Gaussians, the depth regularization $\mathcal{L}_{\text{depth}}$ can be formulated as:
@@
Moreover, we encourage the depth variation of Gaussian primitives to be smooth. We start with adopting the $c$-channel segmentation module from [26] to segment each estimated depth $\mathcal{D}_{i}$ into $c$ similar regions, as shown in part (3) of Fig. 2. For each region, we introduce a local depth constraint $\mathcal{L}_{\text{smooth}}$ to ensure the depth variations are smooth, which are expressed as:
@@
The coarse training objective can be formulated as: $\mathcal{L}_{\text{coarse}}=\mathcal{L}_{\text{color}}+\mathcal{L}_{\text{planar}}$ , where $\mathcal{L}_{\text{color}}$ is the color loss of pixel space using in [11]. $\mathcal{L}_{\text{planar}}$ is expressed as:
@@
III-C Planar Gaussian Refinement via Few-shot Diffusion
To optimize the representation of regions lacking supervision, we introduce the diffusion model [28] to futher refine Gaussian primitives. Note that our approach only utilize few-shot data to fine-tune the diffusion model, i.e., the sparse-view inputs.